Find the series solution of the differential equation

x2(π‘₯+2)π‘¦β€²β€²βˆ’π‘₯𝑦′+(1+π‘₯)𝑦=0.



1
Expert's answer
2021-10-25T03:18:27-0400

Given

"x^{2}(x+2)y''-xy'+(1+x)y=0" this is equation (i)

Let "y=\displaystyle\sum_{n=0}^{\infty}C_nx^n"

Then "y'=\displaystyle\sum_{n=1}^{\infty}C_nnx^{n-1}"

And "y''=\displaystyle\sum_{n=2}^{\infty}C_nn(n-1)x^{n-2}"

Substituting the above into our equation (i)

"x^{2}(x+1)\displaystyle\sum_{n=2}^{\infty}C_nn(n-1)x^{n-2}-x\displaystyle\sum_{n=1}^{\infty}C_nnx^{n-1}"

"+(1+x)\displaystyle\sum_{x=0}^{\infty}C_nx^n=0"

"\displaystyle\sum_{n=2}^{\infty}C_nn(n-1)x^{n+1}-\displaystyle\sum_{n=1}^{\infty}C_nnx^{n}+\displaystyle\sum_{n=0}^{\infty}C_nx^n+\displaystyle\sum_{n=0}^{\infty}C_nx^{n+1}=0"

"\displaystyle\sum_{n=3}^{\infty}C_{n-1}(n-1)(n-2)x^{n}+\displaystyle\sum_{n=1}^{\infty}C_nnx^{n}+\displaystyle\sum_{n=0}^{\infty}C_nx^n"

"+\displaystyle\sum_{n=1}^{\infty}C_{n-1}x^{n}=0"

"\displaystyle\sum_{n=3}^{\infty}C_{n-1}(n-1)(n-2)x^{n}-C_1x-2C_2x^{2}-\displaystyle\sum_{n=3}^{\infty}C_nnx^{n}+C_0"

"+C_1x+C_2x^{2}+\displaystyle\sum_{n=3}^{\infty}C_nx^n+C_0x+C_1x^{2}+\displaystyle\sum_{n=3}^{\infty}C_{n-1}x^n=0"

"C_0+C_0x+(-2C_2+C_2+C_1)x^{2}+[\displaystyle\sum_{n=3}^{\infty}(n-1)(n-2)C_{n-1}"

"-nC_n+C_nx^n+C_{n-1}]x^n=0"

Equating each coefficient by 0

"C_0=0"

"-2C_2+C_2+C_1=0"

"-C_2+C_1=0"

"C_1=C_2"

For "n>2" ,

"(n-1)(n-2)C_{n-1}-nC_n+C_n+C_{n-1}=0"

"[(n-1)(n-2)+1]C_{n-1}+(1-n)C_n=0"

"C_n=\frac{(n-1)(n-2)+1}{(n-1)}C_{n-1}"

"C_n=[(n-2)+\frac{1}{(n-1)}]C_{n-1}"

"C_3=(1+\frac{1}{2})C_2"

"C_3=\frac{3}{2}C_2"

"C_4=(2+\frac{1}{3})C_3"

"C_4=\frac{7}{3}C_3"

"C_5=(3+\frac{1}{4})C_4"

"C_5=\frac{13}{4}C_4"

"C_6=(4+\frac{1}{5})C_5"

"C_6=\frac{21}{5}C_5"

Therefore, the series solution of the differential equation is

"y=\frac{3}{2}C_2x^3+\frac{7}{3}C_3x^4+\frac{13}{4}C_4x^5+\frac{21}{5}C_5x^6+..."